OpenStudy (anonymous):

What is a simpler form of the radical expression? ^3sq root of 125x^21y^24 5x^7|y^8| 15x^21|y^24| 15|x^21|y^24 5x^7y^8

OpenStudy (anonymous):

\[^{3} \sqrt[]{125x^{21}y ^{24^{}}}\]

OpenStudy (nubeer):

ok cube root means power 1/3 so \[125^{\frac{ 1 }{ 3 }} (x ^{21\times \frac{ 1 }{ 3 }})(y ^{24\times \frac{ 1 }{ 3 }})\] can you do from here?

OpenStudy (anonymous):

@nubeer what one is the answer 5x^7|y^8| 15x^21|y^24| 15|x^21|y^24 5x^7y^8